Calculate the surface tension of a liquid which rises 50 cm in a circular tube 0.04 mm in diameter. Relative Density of liquid is 0.8 N/m. Assume the angle of contact is 20°
1
Expert's answer
2020-11-30T14:55:21-0500
"h=\frac{2T\cos{20}}{\rho g r}\\0.5=\frac{2T\cos{20}}{(800)(9.8)(0.02\cdot10^{-3})}\\T=0.042\frac{N}{m}"